OpenStudy (anonymous):

I feel very dumb right now, I forgot how to graph a linear equation when it is only a problem that says this: x-y=2

OpenStudy (dumbcow):

when x=0, y=? when y=0, x =? plot those 2 points and there is your line :)

OpenStudy (ghazi):

substitute one value of x ...find y and plot it on graph ..like x=1,2,3...y=??? and plot it

OpenStudy (anonymous):

So basically choose any number.

OpenStudy (ghazi):

yes

OpenStudy (ghazi):

but only for one variable...like either for x or for y

OpenStudy (anonymous):

Alright. Then I solve for the other variable.

OpenStudy (ghazi):

yes....like if you consider x=1 you'll get y= x-2= 1-2= -1

OpenStudy (ghazi):

now you have one pair (1,-1) similarly find at least one more pair and plot it :)

OpenStudy (anonymous):

I wish math was a lot more simple.

OpenStudy (ghazi):

:) it is ...it's an art...like writing ...i guess art of writing is more complex than maths :)

OpenStudy (anonymous):

I am better on writing than I am math.

OpenStudy (ghazi):

do you write?

OpenStudy (anonymous):

Yes, all the time. The teacher when I went to the class kept some of my work to show future students when she train.

OpenStudy (ghazi):

wow!! :)

OpenStudy (anonymous):

I wish my math was like that.

OpenStudy (ghazi):

you'll have it...with small amount of practice....

OpenStudy (anonymous):

I hope so, I have 28 questions to do all mixed up with various problems but has to do with Linear equations. Then I have a test to do, but I have to get it all done before midnight tonight.

OpenStudy (ghazi):

shoot your questions

OpenStudy (anonymous):

Alright, as for this one I might just do 0 for the x and solve for y. (If I remember the function to do it).

OpenStudy (ghazi):

yes...put x=0 and find y

OpenStudy (anonymous):

0-y=2 would be y=-2.

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

(0,-2) then. Now I just have to graph the thing.

OpenStudy (ghazi):

nope...consider another point let's say x=1 now find y

OpenStudy (ghazi):

what you got?

OpenStudy (anonymous):

y would be -1.

OpenStudy (ghazi):

yes ...now you have (0,-2) and (1,-1) now plot it on graph and join these two points

OpenStudy (anonymous):

Alright, got that one done.

OpenStudy (ghazi):

done...?

OpenStudy (anonymous):

That question lol.

OpenStudy (ghazi):

so no more questions??? :)

OpenStudy (anonymous):

Oh there is more lol, just the one i am working on I am trying to do myself. This one is 2x+3y=-6. It reminds me of the y=mx+b formula.

OpenStudy (ghazi):

so what you have to do in this?

OpenStudy (anonymous):

Graph it, but to do that I will have to find the x and y.

OpenStudy (ghazi):

so do it again...consider x=0,1 and find y at these x